OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs that measure 8cm and 13cm. (Round the answer to the nearest hundreth)

Nnesha (nnesha):

` beg side of right triangle is hypotenuse use formula \[a^2 + b^2 = c^2\]here c is hypotenuse

Nnesha (nnesha):

big**

Nnesha (nnesha):

\[\huge\rm a^2 + b^2 =c^2\] replace variables by their values and solve :-) let me know what you get or if you have any question

OpenStudy (mrnood):

this is the basic equation of the Pythagoras theorem You have been told the length of the 2 short sides, (normally denoted a & b) the eqation for the hypotenuse (normally denoted c) is c^2 = a^2 + b^2 so just put in the numbers fro a & b , and solve the equation to get c

OpenStudy (mrnood):

btw - how did you get to b 'human calculator' if you need to ask this?

OpenStudy (anonymous):

because you get those things for medals @MrNood

Nnesha (nnesha):